OpenStudy (anonymous):

Can someone help me solve? Don't want the answer told to me, want to know how to actually solve it. t^2-t-12/t+1 (t+1/t+3)

jimthompson5910 (jim_thompson5910):

So the problem is \[\Large \frac{t^2-t-12}{t+1} \times \frac{t+1}{t+3}\] right?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

Were you able to factor t^2-t-12 ?

OpenStudy (anonymous):

no.. i have trouble with that

jimthompson5910 (jim_thompson5910):

t^2-t-12 is the same as t^2-1t-12 to factor t^2-1t-12 you need to find two numbers that a) multiply to -12 (the last number) AND b) add to -1 (the middle coefficient)

jimthompson5910 (jim_thompson5910):

are you able to find those two numbers?

OpenStudy (anonymous):

12, -1?

jimthompson5910 (jim_thompson5910):

12 times -1 = -12, so far so good but... 12 plus -1 = 11 we want the two numbers to add up to -1, not 11

OpenStudy (anonymous):

ok, so are we keeping -1 or finding a new set?

jimthompson5910 (jim_thompson5910):

what's another way to multiply to -12?

OpenStudy (anonymous):

(-6,2) (-2,6) (-12,1) (-1,12) am i even on the right track??

jimthompson5910 (jim_thompson5910):

you're getting closer, but you haven't found the right pair yet

jimthompson5910 (jim_thompson5910):

there's another way to multiply to -12

OpenStudy (anonymous):

(-3,4) (-4,3)

jimthompson5910 (jim_thompson5910):

which pair adds to -1 ?

OpenStudy (anonymous):

-4,3

jimthompson5910 (jim_thompson5910):

That means \[\Large t^2-t-12\] factors to \[\Large (t-4)(t+3)\] You can check by expanding out (t-4)(t+3) and you'll get the original expression back.

jimthompson5910 (jim_thompson5910):

So \[\Large \frac{t^2-t-12}{t+1} \times \frac{t+1}{t+3}\] turns into \[\Large \frac{(t-4)(t+3)}{t+1} \times \frac{t+1}{t+3}\] Do you see what to do from here?

OpenStudy (anonymous):

so is it t^3+13/t^2+4 ?

OpenStudy (anonymous):

oops, t^3-11/t^2+4?